OpenStudy (anonymous):

For y= x^2 ln x, find: a) dy b) evaluate dy for x=1, dx= 0.01 Calculus problem, please help!

OpenStudy (anonymous):

product rule for this one \[(fg)'=f'g+g'f\] with \[f(x)=x^2,f'(x)=2x,g(x)=\ln(x),g'(x)=\frac{1}{x}\]

OpenStudy (anonymous):

ooh it is \(dy\) then take that thing and stick a \(dx\) next to it

OpenStudy (anonymous):

Thank you for replying! I'm looking this over so give me a minute. :)

OpenStudy (anonymous):

ok so for part a dy= 2xlnx + x^2/x dx

OpenStudy (anonymous):

and then for part b I just plug in x and dx and that's it?